1. What are the electric field and electric flux of a sphere with a radius of 0.500 m and carries a charGe of +1.00 uc at its center?


1
Expert's answer
2022-03-28T11:02:06-0400

The Gauss's law says

"\phi=\oint EdA=\frac{q}{\epsilon_0}"

"\phi=\frac{1.00*10^{-9}}{8.85*10^{-12}}=113\:\rm N\cdot m^2/C"